# Improving $R(3,k)$ in just two bites

### v2

@@ 3 Preliminaries
Let $G_{2}^{I}$ be the subgraph of $G_{2}$ containing exactly all the edges revealed so far. To reduce notation, let $\mathcal{R}=\mathcal{D}\cap\mathcal{S}\cap\mathcal{M}$ in the following. Let
@@
Suppose $I\subseteq V(G)$ with $|I|=k$. Let $\mathcal{B}_{I}$ be the event that $I$ is independent in $G$. Then, assuming $n$ is sufficiently large,
@@
Let $E_{I,R}$ and $E_{I,B}$ be the sets of pairs not revealed so far in $G_{R}$ and $G_{B}$, and $E_{I}=E_{I,R}\cup E_{I,B}$. The key observation regards how $I$ can be independent at the end. Recall that edges in $O^{+}(I)$ would yield edges in $I$, and thus $\pi_{R}(O^{+}(I))$ and $\pi_{B}(O^{+}(I))$ can not contain edges if $I$ is independent. Thus, if $I$ is independent, all edges in $E_{I}$ must be in $\pi_{R}(C^{+}(I))\cup\pi_{B}(C^{+}(I))$.
@@
$T_{R}=E_{I,R}\setminus\bigcup_{i=1}^{N}{N^{+}(r_{i})\choose 2}$,
$T_{B}=E_{I,B}\setminus\bigcup_{i=1}^{N}{N^{+}(b_{i})\choose 2}$,
@@
which, for sufficiently small $\varepsilon_{1}$ and large enough $k$ implies the result. ∎
@@
Note that since $\varphi$ is chosen uniformly at random, $\varphi(I)$ is a uniformly chosen $k$-subset of $V_{R}\times V_{B}$. For any $\ell_{R},\ell_{B}\in[k]$ with $\ell_{R}\cdot\ell_{B}\geq k$, and writing $\ell_{R}=x_{R}k$, $\ell_{B}=x_{B}k$, we have
@@
$\leq{N\choose\ell_{R}}{N\choose\ell_{B}}\frac{{\ell_{R}\ell_{B}\choose k}}{{N^{2}\choose k}}$
$\leq e^{k}{N\choose\ell_{R}}{N\choose\ell_{B}}\left(\frac{\ell_{R}}{N}\right)^{k}\left(\frac{\ell_{B}}{N}\right)^{k}$
$\leq N^{(x_{R}+x_{B}-2)k}k^{(-x_{R}-x_{B}+2)k}e^{(1+x_{R}+x_{B})k}x_{R}^{(1-x_{R})k}x_{B}^{(1-x_{B})k}$
@@
$\mathbb{P}\left(\bigcup_{\begin{subarray}{c}I\subseteq V(G)\\ |I|=k\end{subarray}}\mathcal{B}_{I}\right)$ $=\mathbb{P}\left(\bigcup_{\begin{subarray}{c}I\subseteq V(G)\\ |I|=k\end{subarray}}\mathcal{B}_{I}\cap\mathcal{R}\right)+\mathbb{P}\left(\bigcup_{\begin{subarray}{c}I\subseteq V(G)\\ |I|=k\end{subarray}}\mathcal{B}_{I}\cap\mathcal{R}^{c}\right)$
$\leq\mathbb{P}\left(\bigcup_{\begin{subarray}{c}I\subseteq V(G)\\ |I|=k\end{subarray}}\mathcal{B}_{I}\cap\mathcal{R}\right)+\mathbb{P}(\mathcal{R}^{c})$
$\leq\sum_{\begin{subarray}{c}I\subseteq V(G)\\ |I|=k\end{subarray}}\mathbb{P}(B_{I}\cap\mathcal{R})+o(1)$
@@
5 The hypergraph Ramsey number $R\left(S_{4}^{(3)},S_{k}^{(3)}\right)$
@@
Therefore, if we want to find a $2$-coloring of the complete hypergraph containing no red $S_{4}^{(3)}$ and no blue $S_{k}^{(3)}$, we are looking for a (red) hypergraph in which no link of any vertex contains a triangle or an independent set of size k1k-1. In other words, we are looking for a hypergraph in which every vertex link is a Ramsey graph for R(3,k1)R(3,k-1). This immediately implies that $R\left(S_{4}^{(3)},S_{k}^{(3)}\right)\leq R(3,k-1)+1$.
Mubayi suggested to us that with only minor modifications, our proof for the lower bound for $R(3,k)$ also gives a lower bound for $R\left(S_{4}^{(3)},S_{k}^{(3)}\right)$, improving the implied constant in a recent paper by Mubayi and Spanier [mubayi2025k4freetriplesystemslarge], which proves a conjecture by Conlon, Fox, He, Mubayi, Suk and Verstraëte [Conlon23] using a nibble argument. We outline the proof below, leaving the details to the reader.
@@
$\left(\frac{1}{2}-o(1)\right)\frac{k^{2}}{\log k}\leq R\left(S_{4}^{(3)},S_{k}^{(3)}\right)\leq\left(1+o(1)\right)\frac{k^{2}}{\log k}$.
@@
    Independent $H(N,p)$ random $3$-graphs $H_{R}$ and $H_{B}$ are generated on vertex sets $V_{R}=\{r_{1},\dots,r_{N}\}$ and $V_{B}=\{b_{1},\dots,b_{N}\}$, where $p=\beta\sqrt{\frac{\log n}{n}}.$ Edges in $H_{R}$ and $H_{B}$ are referred to as red and blue edges respectively.
@@
    The $3$-graph $H_{1}=H_{R}*H_{B}$ on vertex set $V_{R}\times V_{B}$ is constructed. This is similar to the co-normal product of $H_{R}$ and $H_{B}$: $(r,b)(r^{\prime},b^{\prime})(r^{\prime\prime},b^{\prime\prime})\in E(H_{1})$ if at least one of $rr^{\prime}r^{\prime\prime}\in E(H_{R})$ and $bb^{\prime}b^{\prime\prime}\in E(H_{B})$ while $|\{r,r^{\prime},r^{\prime\prime},b,b^{\prime},b^{\prime\prime}\}|=6$ (this last condition differs from the co-normal product, but it simplifies our ensuing analysis). If $rr^{\prime}r^{\prime\prime}\in E(H_{R})$, the resulting edge is colored red. If $bb^{\prime}b^{\prime\prime}\in E(H_{B})$, the resulting edge is colored blue. Edges may, potentially, be colored both red and blue.
@@
    Let $V(H)=\{v_{1},\ldots,v_{n}\}$. Choose an injection $\varphi:V(H)\rightarrowtail V_{R}\times V_{B}$ uniformly at random. The graph induced by $H_{1}$ on $\varphi(V(H))$ is denoted $H_{2}$, which has vertex set $V(H_{2})=\varphi(V(H)).$
  4. 4.
    Create the final $S_{4}^{(3)}$-free graph $H$ as follows:
    • First, red edges in $H_{2}$ are inspected in lexicographical order (i.e., edges are ordered first by their lowest red coordinate, then by their middle coordinate, and then by the highest red coordinate.) If the pre-image (under $\varphi$) of a considered edge does not create a monochromatic red $S_{4}^{(3)}$ with those edges in $H$ previously added, it is added to $H$. Otherwise, it is not.
    • Next, the same is done with blue edges.
    • Then, from any non-monochromatic $S_{4}^{(3)}$ with two blue edges that may remain in $G$, the red edge is removed.
    • Finally, from any non-monochromatic $S_{4}^{(3)}$ with two red edges that may remain in $G$, the blue edge is removed.
This procedure results in an $S_{4}^{(3)}$-free $3$-graph $H$ isomorphic to a subgraph of $H_{2}$ – the first two steps result in no monochromatic $S_{4}^{(3)}$ being created, while the last two steps destroy the remaining copies of $S_{4}^{(3)}$.
The link of any vertex $(v,v^{\prime})\in V(H_{2})$ is exactly the same as the graph $G_{2}$ above (when we replace $n$ by $n+1$, and delete horizontal and vertical edges, which get deleted with high probability when we construct $G$ anyways). So considering pairs $((v,v^{\prime}),I)$ instead of $I$ in $V(G)$ yields the exact same computations as in Section 3. The extra factor of $n$ when counting these pairs can always be swallowed up in the error terms.
The only addition is that some edges in the link of $(v,v^{\prime})$ may be omitted or deleted due to $S_{4}^{(3)}$ in $H_{2}$ not centered at $(v,v^{\prime})$, and these are not captured in our analysis of the sets $X_{w}(I)$ inside the link of $(v,v^{\prime})$. For this to happen, we would need three vertices $x,y,z\in V_{R}$ (or $V_{B}$) with edges $vxy,xyz\in E(H_{R})$ (or $v^{\prime}xy,xyz\in E(H_{B})$) in which case the pair $yz$ would be closed in the link of $(v,v^{\prime})$. For every $k$-set $\varphi(I)\subseteq V(G_{2})$, we expect at most
$\pi_{R}(I)^{2}Np^{2}\leq\frac{k^{2}}{4}\frac{n}{\log^{2}(n)}\frac{\log n}{n}=\frac{k^{2}}{4\log n}$
such closed pairs $yz$ in $\varphi(I)$ coming from $H_{R}$, and the same bound from $H_{B}$. A Chernoff bound then shows that with high probability the number of these pairs is $o(k^{2})$ for all $((v,v^{\prime}),I)$.
Before the proof analogous to Lemma 4.1, we first reveal all edges of $H_{R}$ and $H_{B}$ not containing $v$ and $v^{\prime}$, then we proceed as in the graph case with edges in the link not contained in $\pi_{R}(I)\cup\pi_{B}(I)$, and then iteratively reveal all edges in closed pairs in the link until the process stops.
@@ 6 Conclusion
In any $n$-vertex triangle free graph $G$ we have that $\alpha(G)\geq\Delta(G)\geq\frac{2e(G)}{n}$. Our construction has $\alpha(G)=(1+o(1))\frac{2e(G)}{n}=(1+o(1))\sqrt{n\log n}$; furthermore these independent sets appear both as the neighborhoods of single vertices, as well as ‘mixed’ throughout $G$, resembling those in a random graph of the same density. This is one of the main supporting heuristics for Conjecture 1.1, and we agree that Shearer’s [Shearer83] upper bound on $R(3,k)$ can likely be improved. Improving Shearer’s Theorem is related to the well-studied question of whether there is an efficient algorithm for finding large independent sets in the random graph $G(n,p)$. For most given densities, known efficient algorithms can find only independent sets of about half the size of the independence number.
@@
This gives further support to Conjecture 1.1: if one believes that maximum Ramsey graphs have certain pseudo-random properties, it is unsurprising that a lower bound implied by a bound from a greedy algorithm based on the average degree would be smaller than the actual independence number by a factor of $\sqrt{\frac{1}{2}}$. A conjecture by Davies, Jenssen, Perkins, and Roberts [Davies18] relating maximum and average size of independence sets in triangle-free graphs would also imply the upper bound in Conjecture 1.1.
@@ Appendix A Inequalities
Here we briefly record a few standard inequalities we use, in the notation of Frieze and Karoński [FriezeKarBook].

### v3

@@ 3 Controlling the edge deletions
To reduce notation, let $\mathcal{R}=\mathcal{D}\cap\mathcal{S}\cap\mathcal{M}$ in the following. Let
@@
Suppose $I\subseteq V(G)$ with $|I|=k$. Let $\mathcal{B}_{I}$ be the event that $I$ is independent in $G$. Then, assuming $n>n_{0}$,
@@
Let $E_{I,R}$ and $E_{I,B}$ be the sets of pairs not revealed so far in $G_{R}$ and $G_{B}$, and $E_{I}=E_{I,R}\cup E_{I,B}$. The key observation regards how $I$ can be independent at the end. Recall that if $I$ is independent, all edges in $E_{I}$ must be in $\pi_{R}(C(I))\cup\pi_{B}(C(I))$. More precisely, all edges in $E_{I,R}\cap E(G_{R})$ must be in some $N^{+}(r_{i})\choose 2$} or in some $\pi_{R}\left({N_{b_{i}}\choose 2}\right)$, and all edges in $E_{I,B}\cap E(G_{B})$ must be in some $N^{+}(b_{i})\choose 2$} or in some $\pi_{B}\left({N_{r_{i}}\choose 2}\right)$.
@@
$T_{R}$ $=E_{I,R}\setminus\bigcup_{i=1}^{m}{N^{+}(r_{i})\choose 2}$, $T_{B}$ $=E_{I,B}\setminus\bigcup_{i=1}^{m}{N^{+}(b_{i})\choose 2}$.
@@
which, for sufficiently small $\varepsilon_{1}$ and $k>\sqrt{n_{0}}$ implies the result. ∎
@@
Note that since $\pi$ is chosen uniformly at random, $\pi(I)$ is a uniformly chosen $k$-subset of $V_{R}\times V_{B}$. For any $\ell_{R},\ell_{B}\in[k]$ with $\ell_{R}\cdot\ell_{B}\geq k$, and writing $\ell_{R}=x_{R}k$, $\ell_{B}=x_{B}k$, we have
@@
$\leq{{m}\choose\ell_{R}}{{m}\choose\ell_{B}}\frac{{\ell_{R}\ell_{B}\choose k}}{{{m}^{2}\choose k}}$
$\leq e^{k}{{m}\choose\ell_{R}}{{m}\choose\ell_{B}}\left(\frac{\ell_{R}}{{m}}\right)^{k}\left(\frac{\ell_{B}}{{m}}\right)^{k}$
$\leq{m}^{(x_{R}+x_{B}-2)k}k^{(-x_{R}-x_{B}+2)k}e^{(1+x_{R}+x_{B})k}x_{R}^{(1-x_{R})k}x_{B}^{(1-x_{B})k}$
@@
$\mathbb{P}\left(\bigcup_{I\in{V(G)\choose k}}\mathcal{B}_{I}\right)$ $=\mathbb{P}\left(\bigcup_{I\in{V(G)\choose k}}\mathcal{B}_{I}\cap\mathcal{R}\right)+\mathbb{P}\left(\bigcup_{I\in{V(G)\choose k}}\mathcal{B}_{I}\cap\mathcal{R}^{c}\right)$
$\leq\mathbb{P}\left(\bigcup_{I\in{V(G)\choose k}}\mathcal{B}_{I}\cap\mathcal{R}\right)+\mathbb{P}(\mathcal{R}^{c})$
$\leq\sum_{I\in{V(G)\choose k}}\mathbb{P}(B_{I}\cap\mathcal{R})+o(1)$
@@
5 The hypergraph Ramsey number $R\bigl(S_{4}^{(3)},S_{k}^{(3)}\bigr)$
@@
Therefore, if we want to find a $2$-coloring of the complete hypergraph containing no red $S_{4}^{(3)}$ and no blue $S_{k}^{(3)}$, we are looking for a (red) hypergraph in which no link of any vertex contains a triangle or an independent set of size k1k-1. In other words, we are looking for a hypergraph in which every vertex link is a Ramsey graph for R(3,k1)R(3,k-1). This immediately implies that $R\bigl(S_{4}^{(3)},S_{k}^{(3)}\bigr)\leq R(3,k-1)+1$.
Mubayi suggested to us that with only minor modifications, our proof for the lower bound for $R(3,k)$ also gives a lower bound for $R\bigl(S_{4}^{(3)},S_{k}^{(3)}\bigr)$, improving the implied constant in a recent paper by Mubayi and Spanier [17], which proves a conjecture by Conlon, Fox, He, Mubayi, Suk and Verstraëte [9] using a nibble argument. We outline the proof below, leaving the details to the reader.
@@
$\left(\frac{1}{2}-o(1)\right)\frac{k^{2}}{\log k}\leq R\bigl(S_{4}^{(3)},S_{k}^{(3)}\bigr)\leq\left(1+o(1)\right)\frac{k^{2}}{\log k}$.
@@
    Independent H(m,p)H({m},p) random $3$-graphs $H_{R}$ and $H_{B}$ are generated on vertex sets $V_{R}=\{r_{1},\dots,r_{{m}}\}$ and $V_{B}=\{b_{1},\dots,b_{{m}}\}$, where $p=\beta\sqrt{\frac{\log n}{n}}.$ Edges in $H_{R}$ and $H_{B}$ are referred to as red and blue edges respectively.
@@
    Let $V(H)=\{v_{1},\ldots,v_{n}\}$. Choose an injection $\pi:V(H)\rightarrowtail V_{R}\times V_{B}$ uniformly at random. A triple $v_{i}v_{j}v_{\ell}$ is a red edge in $\tilde{E}_{R}(H)$ if for $(\pi(v_{i}),\pi(v_{j}),\pi(v_{\ell}))=((r,b),(r^{\prime},b^{\prime}),(r^{\prime\prime},b^{\prime\prime}))$, $rr^{\prime}r^{\prime}\in E(H_{R})$ and $b,b^{\prime},b^{\prime\prime}\in V_{B}$ are three different vertices. Blue edges $\tilde{E}_{B}(H)$ are defined symmetrically.
@@
    Create the final $S_{4}^{(3)}$-free (multi)graph $H$ with edge set $E(H)\subseteq\tilde{E}(H)=\tilde{E}_{R}(H)\cup\tilde{E}_{B}(H)$ as follows:
    • For any red copy of $S_{4}^{(3)}$, out of the three red edges in $\tilde{E}_{R}(H)$ considered in lexicographical order (i.e., edges are ordered first by their lowest red coordinate of their image, then by their middle red coordinate, and then by the highest red coordinate), the last edge is deleted.
    • The same is done with blue copies of $S_{4}^{(3)}$, considering the lexicographic order on the blue edges.
    • From any non-monochromatic $S_{4}^{(3)}$ with exactly two blue edges, the red edge is removed.
    • From any non-monochromatic $S_{4}^{(3)}$ with exactly two red edges, the blue edge is removed.
The link of any vertex $v\in V(H)$ is exactly the same as the graph $G$ we constructed for $R(3,k)$. Considering pairs $(v,I)$ instead of $I$ in $V(G)$ yields the exact same computations as in Section 3. The extra factor of $n$ when counting these pairs can always be absorbed in the error terms.
The only addition is that some edges in the link of $v$ may be deleted due to $S_{4}^{(3)}$ in $\tilde{E}(H)$ not centered at $v$, and these are not captured in our analysis of the sets $X_{w}(I)$ inside the link of $v$. For this to happen and $\pi(v)=(v_{R},v_{B})$, we would need three vertices $x,y,z\in V_{R}$ (or $V_{B}$) with edges $v_{R}xy,xyz\in E(H_{R})$ (or $v_{B}xy,xyz\in E(H_{B})$) in which case the pair $yz$ would be closed in the link of $v$. For every $k$-set $\pi(I)\subseteq V_{R}\times V_{B}$, we expect at most
$\pi_{R}(I)^{2}{m}p^{2}\leq\frac{k^{2}}{4}\frac{n}{\log^{2}(n)}\frac{\log n}{n}=\frac{k^{2}}{4\log n}$
such closed pairs $yz$ in $\pi(I)$ coming from $H_{R}$, and the same bound from $H_{B}$. A Chernoff bound then shows that with high probability the number of these pairs is $o(k^{2})$ for all $(v,I)$.
Before the proof analogous to Lemma 4.1, we first reveal all edges of $H_{R}$ and $H_{B}$ not containing $v_{R}$ and $v_{B}$, then we proceed as in the graph case with edges in the link not contained in $\pi_{R}(I)\cup\pi_{B}(I)$, and then iteratively reveal all edges in closed pairs in the link until the process stops.
@@ 6 Conclusion
In any $n$-vertex triangle free graph $G$ we have that $\alpha(G)\geq\Delta(G)\geq\frac{2e(G)}{n}$. Our construction has $\alpha(G)=(1+o(1))\frac{2e(G)}{n}=(1+o(1))\sqrt{n\log n}$; furthermore these independent sets appear both as the neighborhoods of single vertices, as well as ‘mixed’ throughout $G$, resembling those in a random graph of the same density. This is one of the main supporting heuristics for Conjecture 1.1, and we agree that Shearer’s [19] upper bound on $R(3,k)$ can likely be improved. Improving Shearer’s Theorem is related to the well-studied question of whether there is an efficient algorithm for finding large independent sets in the random graph $G(n,p)$. For most given densities, known efficient algorithms can find only independent sets of about half the size of the independence number.
@@
This gives further support to Conjecture 1.1: if one believes that maximum Ramsey graphs have certain pseudorandom properties, it is unsurprising that a lower bound implied by a bound from a greedy algorithm based on the average degree would be smaller than the actual independence number by a factor of $\sqrt{\frac{1}{2}}$.
@@ Appendix A Inequalities
Here we briefly record a few standard inequalities we use, in the notation of Frieze and Karoński [14].